### v3

# Rethinking Open-Vocabulary Segmentation of Radiance Fields in 3D Space

Abstract
Understanding the 3D semantics of a scene is a fundamental problem for various scenarios such as embodied agents. While NeRFs and 3DGS excel at novel-view synthesis, previous methods for understanding their semantics have been limited to incomplete 3D understanding: their segmentation results are rendered as 2D masks that do not represent the entire 3D space. To address this limitation, we redefine the problem to segment the 3D volume and propose the following methods for better 3D understanding. We directly supervise the 3D points to train the language embedding field, unlike previous methods that anchor supervision at 2D pixels. We transfer the learned language field to 3DGS, achieving the first real-time rendering speed without sacrificing training time or accuracy. Lastly, we introduce a 3D querying and evaluation protocol for assessing the reconstructed geometry and semantics together. Code, checkpoints, and annotations are available at the project page.
Project page — https://hyunji12.github.io/Open3DRF/
1 Introduction
Semantically understanding 3D space is important for various computer vision tasks. For instance, it is crucial to segment 3D objects accurately for robot manipulation (Rashid et al. 2023; Zheng et al. 2024). Recently, several works have focused on understanding 3D scenes represented by radiance fields such as Neural Radiance Fields (NeRFs) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). LERF (Kerr et al. 2023) introduces a language embedding field which is rendered on a chosen viewpoint to be queried by open vocabulary. The language field is supervised by CLIP (Radford et al. 2021) embeddings from the multi-scale patches to capture various sizes of objects. Subsequent works (Zhang, Li, and Ahuja 2024; Qin et al. 2023) incorporate SAM masks (Kirillov et al. 2023) to supervise the language field for clear segmentation boundaries.
We revisit the 3D understanding of NeRFs and 3DGS in four aspects: problem setting, supervision, embeddings, and evaluation. The problem setting of previous works leads to limited 3D understanding as they merely produce rendered 2D masks for given viewpoints rather than 3D semantic volumes. On the other hand, we set the problem to segment 3D volume regarding the semantics of 3D points.
The previous methods encourage the rendered embeddings, rather than the embeddings in 3D space, to match the ground truth. Furthermore, their multi-scale language embeddings require finding the optimal scale for a given viewpoint and a query text, which is prone to multi-view inconsistency. In contrast, we encourage 3D points to learn language embeddings following our revisited problem setting. Although the ground truth is still language embeddings of 2D images, changing the dimension for computing the loss greatly improves 3D understanding as shown in Figure 1. In addition, we remove multi-scale embeddings by encoding masked objects, achieving multi-view consistent rendered feature maps.
Meanwhile, previous works on understanding 3DGS explicitly add language embeddings for each Gaussian and jointly train them (Qin et al. 2023; Zhou et al. 2023). However, directly appending a 512-dimensional language embedding to each Gaussian and rendering them result in out-of-memory (OOM) issues. Hence, they either 1) compress language embeddings to low-dimensional features, significantly sacrificing the accuracy (Shi et al. 2023; Qin et al. 2023), or 2) modify the rasterizer, which is slow for training (Zhou et al. 2023; Ye and Kanazawa 2023). To overcome these drawbacks, we transfer our learned language field into 3DGS to enable the real-time rendering of the accurate language field.
Our final aspect is evaluation. Existing works measure mIoU in pixels between the rendered 2D masks and the 2D ground truth, or measure mIoU on ground truth point clouds ignoring the reconstructed geometry (Engelmann et al. 2024). However, 3D understanding should know the correct 3D volume of a target semantics. Therefore, the evaluation method should assess semantics and the reconstructed geometry together. Accordingly, we propose to evaluate the accuracy of 3D understanding as the agreement between estimated volume in mesh and ground truth mesh, measured in F1-score. The proposed evaluation is applicable to both NeRF and 3DGS.
In the experiments, we demonstrate the superiority of our method regarding 1) 3D and rendered 2D segmentation accuracy, 2) training and rendering time, and 3) consistency across viewpoints.
In summary, our contributions are:
We propose a practical problem setting for 3D understanding of NeRFs and 3DGS.
We propose to directly supervise 3D points before volume rendering to learn a language embedding field. It achieves the state-of-the-art accuracy in 3D and rendered 2D segmentation.
We propose to transfer the language field to 3DGS. It achieves the first real-time rendering speed among open-vocabulary methods which is 28 faster than the previous fastest method.
We propose a 3D evaluation protocol between estimated volume and ground truth volume represented as meshes.
2 Related Work
In this section, we briefly review neural 3D scene representations, especially NeRF and 3DGS, and then discuss the semantic understanding of 3D scenes.
Radiance Fields
NeRF (Mildenhall et al. 2020) reconstructs a scene as a continuous function that maps 3D points to radiance and density values. Volume rendering pipeline renders the points to determine the color of each pixel on an image from an arbitrary perspective. Rather than volumetric rendering, 3DGS (Kerbl et al. 2023) achieves fast rendering speeds by projecting 3D Gaussians onto the camera plane followed by depth-sorted alpha-blending. Each 3D Gaussian stores location, rotation, scale, opacity, and color. Recent studies use these representations to semantically understand 3D scenes for various applications such as robotics (Rashid et al. 2023; Wang et al. 2023; Zheng et al. 2024; van Oort et al. 2024; Li and Pathak 2024).
Semantic Understanding of Radiance Field
The most straightforward way to understand 3D scenes represented by neural radiance fields is by adding an auxiliary branch for semantic segmentation (Zhi et al. 2021; Bing, Chen, and Yang 2022; Liu et al. 2023; Ye et al. 2023). This approach allows synthesizing semantic masks from novel views but requires a pre-defined list of target classes before training, called closed-set. Therefore, retraining or using additional models for untrained queries is necessary, limiting application for open-vocabulary scenarios. As vision-language model (VLM) features (e.g., CLIP (Radford et al. 2021)) expand the semantic understanding to open-set, it becomes a widely used approach to distill CLIP features into 3D scenes.
LERF (Kerr et al. 2023) pre-computes multi-scale patches to prepare multi-scale ground-truth CLIP features. Similar to LERF, FMGS (Zuo et al. 2024) pre-computes multi-scale ground-truth CLIP features and averages them to generate a low-resolution hybrid feature map for training. Instead of using patches, LEGaussians (Shi et al. 2023) distills pixel-level CLIP features. Similarly, OpenNeRF (Engelmann et al. 2024) computes pixel-level CLIP features using OpenSeg (Ghiasi et al. 2022). GOI (Qu et al. 2024) trains the codebook to compress the high-dimensional semantic features into low-dimensional vectors111At the time of this work, the code of GOI was unavailable. However, its released code showed 15 fps on LERF dataset, falling short of real-time rendering.. As it produces mixed CLIP features to contain multiple objects into a patch, LangSplat (Qin et al. 2023) utilizes SAM (Kirillov et al. 2023) to create patches for a single object. OpenMask3D (Takmaz et al. 2023) creates a bounding box from the SAM mask, and crops the image to make per-mask features. However, OpenMask3D uses bounding box, which still leads to mixed CLIP features. Since studies using multi-scale feature maps select the most relevant scale, different scales can be chosen at different views with the same query, leading to multi-view inconsistent results (Kerr et al. 2023; Qin et al. 2023; Shi et al. 2023). On the other hand, we use SAM to make the field free from scales.
Language Embedded 3DGS
Directly embedding high-dimensional CLIP features into each Gaussian is infeasible due to the limited GPU shared memory. Therefore, recent studies for understanding open-vocabulary 3D segmentation using 3D as a backbone address this problem in two ways. LEGaussians and LangSplat compress the CLIP features using an autoencoder (Shi et al. 2023; Qin et al. 2023). However, they need to optimize the autoencoder for each scene and endure performance degradation. Others using 3DGS without feature compression (Zhou et al. 2023; Labe et al. 2024) utilize global memory, suffering for longer training time. Meanwhile, we effectively optimize 3DGS transferred from our learned language field, without the above trade-off.
3 Methods
First, we provide preliminary on LERF (Kerr et al. 2023). Then we redefine 3D segmentation of radiance fields as computing 3D relevancy scores. Accordingly, we introduce point-wise semantic loss to supervise ray points in 3D space by our scale-free embeddings. In addition, we propose to transfer our learned language field into 3DGS for real-time rendering. Lastly, we introduce the evaluation protocol for 3D segmentation. Figure 2 illustrates an overview.
3.1 Preliminary: LERF
For open-vocabulary segmentation, LERF builds an additional language field on top of iNGP (Müller et al. 2022). The language field is jointly trained with the radiance field by querying point embeddings at sample points’ position and scale along each ray. The 3D point language embeddings are then accumulated via the volume rendering (Max 1995) to obtain rendered language embedding in 2D pixel space: . The weight of each sampled point is calculated as , where is transmittance, is the distance between adjacent samples, and is the volume density. Then, the rendered embedding is normalized to the unit sphere as in CLIP: To train the language field, LERF crops the training dataset into multi-scale patches, creating ground truth language embedding and maximizing the cosine similarity between the rendered language embedding :
Furthermore, LERF builds an additional branch for the DINO (Caron et al. 2021) feature field as an extra regularizer for achieving clearer object boundaries. Similar to the above, the DINO branch is trained to maximize the similarity between rendered DINO embedding and the DINO ground truth . The DINO branch is not used during inference.
For rendered 2D segmentation, LERF computes the relevancy score using the rendered embeddings and obtains the 2D mask:
3.2 Task Redefinition
Existing methods provide limited 3D semantic understanding for given texts via rendered 2D masks on viewpoints that do not directly represent the entire 3D space. Instead, we propose 3D segmentation as a more practical problem setting: segmenting the 3D volume for given texts. While following the basic elements of language embeddings and relevancy scores, we compute relevancy scores of the language embeddings queried on 3D points instead of the ones rendered on 2D images:
where represents predefined canonical texts such as photo and image. In NeRFs, we compute the relevancy scores on the points along the rays through pixels. For 3DGS, we compute the relevancy scores on the center positions of the Gaussians. The regions where the computed relevancy scores surpass the selected threshold are considered object regions.
3.3 Supervising Semantics in 3D Space
It is a reasonable choice to minimize the error between the rendered colors and the ground truth colors on 2D images for novel view synthesis. In contrast, the similar objective for language embeddings (Eq. (1)) harms correctly understanding 3D semantics as shown in Figure 1.
To address this issue, we propose a point-wise semantic loss which uses CLIP embeddings as direct ground truth for the embeddings of 3D points on the ray . Specifically, we maximize the similarity between point embeddings and the ground truth CLIP embedding :
We also apply the same approach described above for DINO regularization. We show that point-wise semantic loss improves not only 3D segmentation but also rendered 2D segmentation thanks to a better understanding of 3D semantics.
Notably, multi-scale approaches (Kerr et al. 2023; Qin et al. 2023) lead to multi-view inconsistency because these approaches have language embedding per scale and different views may have different optimal scales. To address this, we use SAM (Kirillov et al. 2023) to obtain object masks and create scale-free ground truth CLIP embeddings by masking out their background and cropping tightly to the masks. It ensures view-consistent language fields without the need to determine the optimal scale. We note that previous methods with SAM still model multi-scale CLIP embedding fields (Zhang, Li, and Ahuja 2024; Qin et al. 2023).
3.4 Transferring Language Field into 3DGS
A straightforward approach for the same task with 3DGS is to jointly optimize the Gaussians and their additional language embeddings, which requires vast memory consumption. It suffers from slow training222Training on GPU global memory is slow due to the gradient computation in the backward pass. The forward pass is still fast. (Zhou et al. 2023), or requires an additional model for feature compression, which degrades accuracy due to compression loss (Shi et al. 2023; Qin et al. 2023). Appendix provides more details.
To tackle this problem, we propose to simply bake our learned language field into 3DGS by querying the language embedding at the center of the Gaussians. The queried embeddings are frozen. It runs instantly (22ms) and still allows fast rendering22footnotemark: 2.
In addition, we propose to initialize the center coordinates of 3DGS from the learned radiance field. We collect ray points and extract the top 1M points regarding density following NeRFstudio (Tancik et al. 2023). Unlike previous methods (Niemeyer et al. 2024), we implement this approach to align the geometry of 3DGS with the language field. Therefore, we freeze the center position of the Gaussians , without densification or pruning during training. We then transfer language embeddings at the center of Gaussians.
Similar to previous methods, we modify the rasterizer to render high-dimensional language features (see Appendix). The rendered language embedding is obtained by -blending:
where denotes the number of Gaussians overlapping on the pixel. The alpha value is calculated as where denotes the Gaussian kernel and denotes the opacity of Gaussians. We use Eq. (3) for 3D segmentation.
3.5 3D Semantic Evaluation Protocol
A previous work (Engelmann et al. 2024) measures the mIoU between the prediction and annotation on the ground truth point cloud. It might incorrectly classify the points with zero density and a correct embedding, shown in Figure 5-(a).
To address this problem, we propose to measure the F1-score between the exported mesh from the segmented results to ground truth mesh, inspired by surface reconstruction literature (Knapitsch et al. 2017; Li et al. 2023). Precision computes the ratio of correct volume among the estimated volume. Recall computes the ratio of covered volume among the GT volume. As it is impractical to compute the volume, we approximate the volume by a regular grid points in the estimated and GT meshes. A point is considered correct or covered if there exists a point within a radius in the counterpart. We note that this protocol can be generally applied to various neural representations, such as NeRF or 3DGS.
4 Experiments
In this section, we evaluate our methods on various datasets and compare them to competitors in terms of 3D and rendered 2D segmentation with given text queries. We choose LERF, OpenNeRF, LEGaussians, and LangSplat as competitors, which are open-sourced. We re-evalute the competitors using the official code333The official implementation of LangSplat includes evaluation views during training, while we exclude them for fairness, following LERF and OpenNeRF.. Appendix provides details of the datasets.
Backbone | Method | Replica |
F1-score | ||
NeRF | LERF | 0.0845 |
OpenNeRF | 0.0361 | |
Ours-NeRF | 0.1520 | |
3DGS | LEGaussians | 0.0067 |
LangSplat | 0.0087 | |
Ours-3DGS | 0.1353 |
Evaluation
For quantitative comparison in 3D segmentation, we export meshes using the Poisson surface reconstruction. In LERF, OpenNeRF, and Ours-NeRF, we use Nerfstudio (Tancik et al. 2023) to export meshes with 30K sampled points. In LEGaussians, LangSplat and Ours-3DGS, we use SuGaR (Guédon and Lepetit 2023) to export meshes. For rendered 2D segmentation, we evaluate the mIoU and mAP between the ground truth mask and the predicted mask. For qualitative comparison in 3D segmentation, we use 50K points for Nerfstudio in LERF, OpenNeRF, and Ours-NeRF.
4.1 Segmentation Accuracy
3D Segmentation
We compare qualitative results of 3D segmentation by exporting meshes for the region obtained through 3D querying (Section 3.2). In Table 1, both of our models surpass the competitors.
In Figure 6, unlike competitors, Ours-NeRF and Ours-3DGS produce clear segmentation boundaries of the target object from the 3D scene. Notably, both of our models accurately segment complex shape objects and multiple objects like porcelain hand and rubber ducks. In the rubber duck query, LERF, LEGaussians, and LangSplat completely fail to localize the target object, while OpenNeRF incorrectly segments unrelated objects (e.g., a toy elephant and Jake) along with the target object.
Method | LERF | 3D-OVS | ||
mIoU | mAP | mIoU | mAP | |
LERF | 31.88 | 30.44 | 52.60 | 57.03 |
OpenNeRF | 26.52 | 25.84 | 75.12 | 75.35 |
Ours-NeRF | 46.37 | 45.86 | 77.46 | 84.65 |
LEGaussians | 21.43 | 20.81 | 47.98 | 50.86 |
LangSplat33footnotemark: 3 | 37.53 | 36.39 | 74.54 | 79.49 |
Ours-3DGS | 44.37 | 44.57 | 77.51 | 84.86 |
Rendered 2D Segmentation
In Table 2, both of our models show the highest rendered 2D segmentation performance on LERF and 3D-OVS datasets. Also, we present qualitative results of rendered 2D segmentation by comparing a heatmap for given text queries through 2D querying (Section 3.2). As shown in Figure 7, Ours-NeRF and Ours-3DGS show clear segmentation boundaries. LangSplat and LEGaussians occasionally fail to find the target object, as seen with waldo.
Figure 8 shows relevancy maps and optimal scales along different viewpoints. The float value on the top right of the image represents the optimal scale for LERF and LangSplat from each viewpoint. In Figure 8, our method renders view consistent relevancy maps with napkin query. However, LERF and LangSplat show view inconsistent segmentation results due to the changing optimal scale.
Backbone | Method | Training | Rendering | ||||
Render | Decode | Post-Process | FPS | Real-Time | |||
NeRF | LERF | 40 mins | 23688. ms | - | - | 0.04 | ✗ |
OpenNeRF | 40 mins | 5944.7 ms | - | - | 0.17 | ✗ | |
Ours-NeRF | 40 mins | 2337.7 ms | - | - | 0.43 | ✗ | |
3DGS | LEGaussians | 90 mins | 15.665 ms | 384.6 ms | - | 2.50 | ✗ |
LangSplat | 100 mins | 17.249 ms | 0.935 ms | 10473 ms | 0.10 | ✗ | |
Ours-3DGS | 40+5 mins | 14.257 ms | - | - | 70.1 | ✓ |
Method | Replica | LERF | Computational Cost | |||
F1-score | mIoU | mAP | Training Time | FPS | # of Gaussians | |
Ours-3DGS | 0.1354 | 44.37 | 44.57 | 5 mins | 70.1 | 1M |
w/o NeRF init | 0.1108 | 36.14 | 36.74 | 13 mins | 41.2 | 2M |
4.2 Computational Time
Table 3 shows the computational times of the LERF waldo_kitchen scene with a single RTX A5000. We report the training time and rendering time for both ours and the competitors. LangSplat and LEGaussians train an autoencoder for each scene and require decoding during rendering. Note that LangSplat trains 3DGS from scratch, optimizes the language-embedded 3DGS per scale, heavily depends on post-processing (see Appendix). Ours-3DGS takes 40 minutes to train Ours-NeRF, 5 minutes to optimize 3DGS, and 22 milliseconds to query and transfer language field into 3DGS. Ours-3DGS achieves real-time rendering of language embeddings for the first time, 28 faster than LEGaussians.
Method | Replica | LERF | |
F1-score | mIoU | mAP | |
Ours full | 0.1532 | 46.37 | 45.86 |
w/o | 0.0537 | 44.50 | 44.05 |
w/o | 0.1114 | 30.03 | 28.85 |
4.3 Ablation Study
Point-wise Semantic Loss
We demonstrate the necessity of our point-wise semantic loss. Figure 9 qualitatively shows improvements due to point-wise semantic loss compared to 2D supervision with Eq. (1). Point-wise semantic loss removes meaningless floaters related to depth ambiguity. Table 5 shows that leveraging point-wise semantic loss instead of Eq. (1) makes better understanding in 3D space, improving the F1-score (+0.0995). Furthermore, point-wise semantic loss also indirectly improves rendered 2D segmentation accuracy.
Scale-Free Embedding
Our scale-free language embedding field from SAM-segmented objects produce greatly improved accuracy with clear object boundaries compared to multi-scale embedding field. Figure 9 shows that scale-free embedding helps covering the head of the toy elephant which was lost with multi-scale embedding. In Table 5, the quantitative performance of the 3D and rendered 2D segmentation greatly improves using scale-free embedding.
NeRF Initialization
We demonstrate the necessity of initializing 3DGS with the extracted point clouds from our learned NeRF. As shown in Table 4, training 3DGS with SfM (Schönberger and Frahm 2016) initialization with the densification and pruning leads to performance degradation in both 3D and rendered 2D segmentation. This indicates that aligning geometry and semantics is essential for transferring the language field to 3DGS. Moreover, NeRF initialization produces fewer Gaussians (0.50) which lead to less training time (0.38) and faster rendering speed (1.70).
5 Conclusion
We revisit the current literature on 3D understanding of NeRF and 3DGS and revise the problem setting. We reformulate the task to produce 3D segmented volumes instead of rendered 2D masks and propose a 3D evaluation protocol. We achieve state-of-the-art segmentation accuracy in both 3D and rendered 2D by computing loss directly on 3D points. Moreover, we enable the first real-time rendering speed among open-vocabulary methods by transferring the learned language field to 3DGS. We hope this paper drives forward a better 3D understanding of radiance fields by reconsidering the problem set.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00439762, Developing Techniques for Analyzing and Assessing Vulnerabilities, and Tools for Confidentiality Evaluation in Generative AI Models)
Appendix A Experiments Details
A.1 Implementation Details
We train the NeRF and the CLIP branch jointly from scratch with the same configurations of LERF except the learning rate on Replica datasets which is 1e-4. The CLIP features are extracted using OpenCLIP ViT-B/16 trained on LAION-2B. We use an automatic mask generator from ViT-H SAM with the default configuration.
We optimize 3DGS in Section 3.4 with the same settings as the official 3DGS, except for the following modifications: 1) freezing the position of the Gaussian center , 2) setting the rotation learning rate to 1e-4, and 3) running for 10K iterations without a densification process.
We train our models and competitors with a single RTX A5000.
A.2 Details of Datasets
We use three datasets for evaluation: Replica dataset for 3D segmentation, and LERF and 3D-OVS datasets for rendered 2D segmentation. All of these datasets are widely used to evaluate the segmentation performance of radiance fields. We fix the thresholds for each dataset as detailed in Table 9 to evaluate the performance.
Replica dataset.
Replica dataset is a high-quality indoor 3D mesh dataset with semantic labels. We measure 3D segmentation performance with labeled 3D mesh, using class names as the text queries.
LERF dataset.
LERF dataset is a collection of real-world scenes captured by iPhone. This includes text queries and their ground truth 2D bounding boxes. In LERF dataset, most competitors (Qin et al. 2023; Shi et al. 2023) report their performance by building their own ground truth mask datasets corresponding to their partly chosen text queries. However, it is a challenge to compare because the performance of open-vocabulary segmentation can significantly vary depending on the text query.
Hence, starting from the given text queries and 2D bounding boxes, we use SAM to create auxiliary masks and manually correct errors. SAM(Kirillov et al. 2023) receives an image and a user prompt (bounding box or point) as input and creates three masks. Among these candidate masks, we select the ground truth mask that corresponds most closely to the queried object. Then, we manually label a few inaccurate masks. The masks will be publicly available. We use four scenes figurines, ramen, teatime, waldo_kitchen from LERF dataset. We choose three evaluation views that are not blurry. These masks will be publicly available online.
3D-OVS dataset.
3D-OVS dataset is a collection of real-world scenes, text queries, and their ground truth pixel-level labels. The dataset consists of 10 scenes, each with 20-30 viewpoints, including five evaluation viewpoints. Several scenes in 3D-OVS dataset show poor reconstruction quality from the evaluation view shown in Figure 5. Therefore, we report the performance only for the scenes with successful reconstruction: (bed, bench, lawn, and office_desk).
ScanNet dataset.
ScanNet dataset is a real-world dataset that includes RGB-D scans, meshes, and instance-level semantic labels. It is widely used to evaluate performance in 3D scene understanding tasks. In our case, we select one scene (scene0003) and provide the 3D segmentation accuracy, in Appendix C.
Appendix B Details of Previous Language Embed 3DGS
B.1 Splatting High Dimensions Embeddings on 3DGS
To jointly train 3DGS with language features, a straightforward approach is to embed the 512-dimensional language feature into each Gaussian. However, this approach results in out-of-memory(OOM) issues. Splatting 512-dimensional language embeddings requires 522 KB of shared memory per streaming multiprocessor (SM). Specifically, 3DGS (Kerbl et al. 2023) requires 10 KB of shared memory and needs an additional 1 KB/SM for each increase in the dimension of the language feature. 3DGS allocates static shared memory, and CUDA limits this allocation to 48 KB for architectural compatibility. Even with a dynamic allocation that enables allocation above the 48 KB limit, 522 KB of shared memory exceeds the maximum shared memory of current top GPUs (e.g., the H100 has a maximum shared memory of 228 KB).
Consequently, existing methods address these challenges using the following strategies, each with its trade-offs:
Allocate rasterizer variables for the backward pass to global memory instead of shared memory (Zhou et al. 2023). While this avoids performance degradation, it significantly slows down training.
Use dynamic shared memory allocation with reduced tile size in the rasterizer (Ye et al. 2024). This method is faster than global memory allocation while using more shared memory, but it still slows down training due to the smaller tile size.
Render in chunks (Bhalgat et al. 2024). Although this accelerates training and prevents performance degradation, it drops fps (10 fps reported in the original paper).
Compressing the features produces poor mIoU (-7.30) but speeds up the training time 47, as shown in Table 11. However, Ours-3DGS achieves fast training without feature compression.
Method | mIoU | mAP | Training Time |
3dim-feature | 31.54 | 30.09 | 30 mins |
512dim-feature | 38.84 | 38.20 | 1400 mins |
B.2 Post-process of LangSplat
We show the performance difference with and without post-processing in LangSplat, as shown in Figure 6. Without post-processing, it achieves a fast rendering speed of 55 fps (550) 444The post-processing consists of a mean convolution filter on the relevancy map implemented with OpenCV (61.98ms) and a mean convolution filter on the binary mask implemented with nested for-loops (3249 ms). This is done for three scales, 10473 ms in total. but produces a poor mIoU (-5.5) as shown in Table 12. Note that Ours-3DGS achieve faster rendering speed even compared to LangSplat without post-processing, as shown in Table 3.
LangSplat | mIoU | mAP |
w/ post-process | 37.53 | 36.39 |
w/o post-process | 32.01 | 30.64 |
Appendix C More Results
We mainly evaluate 3D segmentation accuracy on Replica datasets. To expand this, we also evaluate and compare on scene_0003 of ScanNet (Dai et al. 2017) dataset, in Table 13.
Additionally, we report the quantitative results of each scene on Replica dataset, LERF dataset, and 3D-OVS dataset, in Table 14-16.
Backbone | Method | ScanNet |
F1-score | ||
NeRF | LERF | 0.0333 |
OpenNeRF | 0.0663 | |
Ours-NeRF | 0.0849 | |
3DGS | LEGaussians | 0.0064 |
LangSplat | 0.0340 | |
Ours-3DGS | 0.0769 |
Backbone | Method | office_0 | office_1 | office_2 | office_3 |
F1-score | F1-score | F1-score | F1-score | ||
NeRF | LERF | 0.0774 | 0.0191 | 0.0644 | 0.0621 |
OpenNeRF | 0.0486 | 0.0169 | 0.0318 | 0.0356 | |
Ours-NeRF | 0.1434 | 0.0434 | 0.1493 | 0.1255 | |
3DGS | LEGaussians | 0.0013 | 0.0043 | 0.0109 | 0.0031 |
LangSplat | 0.0004 | 0.0123 | 0.0044 | 0.0038 | |
Ours-3DGS | 0.1097 | 0.0406 | 0.1395 | 0.1181 | |
Backbone | Method | office_4 | room_0 | room_1 | room_2 |
F1-score | F1-score | F1-score | F1-score | ||
NeRF | LERF | 0.0645 | 0.1124 | 0.1607 | 0.1158 |
OpenNeRF | 0.0521 | 0.0180 | 0.0432 | 0.0423 | |
Ours-NeRF | 0.1722 | 0.1654 | 0.1934 | 0.2232 | |
3DGS | LEGaussians | 0.0111 | 0.0080 | 0.0066 | 0.0085 |
LangSplat | 0.0107 | 0.0139 | 0.0122 | 0.0118 | |
Ours-3DGS | 0.1629 | 0.1636 | 0.1529 | 0.1959 |
Method | figurines | ramen | teatime | waldo_kitchen | ||||
mIoU | mAP | mIoU | mAP | mIoU | mAP | mIoU | mAP | |
LERF | 40.67 | 38.75 | 24.56 | 24.35 | 36.85 | 35.06 | 25.44 | 23.62 |
OpenNeRF | 16.24 | 16.02 | 23.05 | 23.57 | 39.08 | 38.26 | 27.69 | 25.50 |
Ours-NeRF | 59.79 | 59.00 | 37.44 | 38.38 | 47.59 | 46.92 | 40.67 | 39.16 |
LEGaussians | 21.02 | 20.57 | 24.90 | 24.67 | 24.50 | 24.09 | 15.28 | 13.89 |
LangSplat | 42.23 | 40.55 | 48.17 | 48.71 | 37.74 | 36.26 | 21.98 | 20.05 |
Ours-3DGS | 56.23 | 55.79 | 38.28 | 40.81 | 41.59 | 41.16 | 42.60 | 41.79 |
Method | bed | bench | lawn | office_desk | ||||
mIoU | mAP | mIoU | mAP | mIoU | mAP | mIoU | mAP | |
LERF | 45.29 | 57.74 | 50.88 | 51.59 | 54.19 | 57.40 | 60.04 | 61.41 |
OpenNeRF | 82.25 | 82.40 | 70.24 | 71.14 | 78.70 | 79.27 | 69.29 | 69.79 |
Ours-NeRF | 66.71 | 78.65 | 66.69 | 78.81 | 96.74 | 96.98 | 79.72 | 84.18 |
LEGaussians | 33.23 | 44.94 | 63.86 | 64.25 | 43.48 | 44.20 | 51.36 | 50.05 |
LangSplat | 65.74 | 73.01 | 82.75 | 88.06 | 82.41 | 87.95 | 67.26 | 68.93 |
Ours-3DGS | 67.40 | 79.33 | 67.08 | 79.20 | 96.25 | 96.92 | 78.89 | 84.36 |